[figure: 9.7 cm]

EVIDENCE FOR A TEMPERATURE AND FIELD DEPENDENCE OF THE ANISOTROPY OF SUPERCONDUCTING MgB_{2}

R. Puzniak^{a}, M. Angst^{b}, A. Wisniewski^{a}, J. Jun^{b}, S.M. Kazakov^{b}, J. Karpinski^{b}, J. Roos^{c}, H. Keller^{c} aInstitute of Physics, PAS, Al. Lotników 32/46, PL-02-668 Warsaw, Poland bSolid State Physics Laboratory ETH, CH-8093 Zurich, Switzerland cInstitute of Physics, University of Zurich, CH-8057 Zurich, Switzerland

The anisotropy γ of superconducting state of high quality single-crystals of MgB_{2} was determined using torque magnetometry. The anisotropy of the upper critical field decreases with increasing temperature from γ ≈ 6 at 15 K to 2.8 at 35 K. Measurements of the reversible torque near T_{c} reveal an almost linear field dependence of the anisotropy of the coherence length and/or the penetration depth as well. The anisotropy increases nearly linearly from γ ≈ 2 in zero field to 3.7 in 1 T. Our results imply a breakdown of standard anisotropic Ginzburg-Landau theory with a temperature and field independent effective mass anisotropy. The unusual temperature dependence of γ is a true bulk property and can be explained by non-local effects of anisotropic pairing and/or the wave-vector dependence of the effective mass tensor.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : R. Puzniak

Address for correspondence : Instytut Fizyki, PAN Al. Lotników 32/46, 02-668 Warszawa

Email address : puzni@ifpan.edu.pl